SIAS Web Interface
SIAS Date: 2017-05-25Time of Printout: 00:33
May 2017
SunMonTueWedThuFriSat
 
25
 
To ->.
Go to
Telescope
Status
K1-NOT READY K2-NOT READY
Instrument
Status
LRIS-ADC NOT READY
lris-adc(6)
DEIMOS NOT READY
deimos(19)
ObserversS. KulkarniHennawi
SAPercy G.Percy G.
OACynthia W.
rem-ops
Gary P.
summit
SWOCSylvain C.
Who System Description Last Changed Job Done Comments
Initials Time
SupportLRIS**HIGH PRIORITY** Fill LN after arriving in the morning.
2010-03-08 17:03
   
SupportESI**HIGH PRIORITY** Fill LN on arrival
2010-02-15 12:02
   
SupportLRIS
LRIS Configuration:

SLITMASK1: direct BC=5920-no change
SLITMASK2: GOH_LRIS BC=2466-no change
SLITMASK3: long_0.7 BC=5115
SLITMASK4: long_1.5 BC=5117
SLITMASK5: long_1.0 BC=5116-no change
SLITMASK6: gws_L3b. BC=3662-no change
SLITMASK7: gws_L2.f BC=3659-no change
SLITMASK8: dsf_L2.f BC=3664-no change
SLITMASK9: dsfbck.f BC=3663-no change
SLITMASK10: m2135L.f BC=3649-no change
GRATING1: mirror-no change
GRATING2: 400/8500
GRATING3: Blank (Bad port)-no change
GRATING4: 600/7500
GRATING5: 831/8200-no change
RED_FILT1: Clear-no change
RED_FILT2: R-no change
RED_FILT3: RG850
RED_FILT4: I-no change
RED_FILT5: OG570
RED_FILT6: GG495-no change

No changes required for any
position in the following:
GRISM, Blue Filter, DICHROIC
2017-05-24 08:05
 00:00 
SupportDEIMOS
DEIMOS Configuration:

SLITMASK2: m1143 BC=10727
SLITMASK3: m231 BC=10714
SLITMASK4: m159 BC=10712
SLITMASK5: m183 BC=10713
SLITMASK6: LVMslitC BC=9400-no change
SLITMASK7: m239 BC=10715
SLITMASK8: m247 BC=10716
SLITMASK9: m323 BC=10717
SLITMASK10: m338 BC=10718
SLITMASK11: m1335 BC=10729
SLITMASK12: GOH_X BC=10160-no change
SLITMASK13: Empty_13 (Bad port)-no change
FILTER Pos 1: V
FILTER Pos 2: Z
FILTER Pos 3: OG550-no change
FILTER Pos 4: I-no change
FILTER Pos 5: GG495-no change
FILTER Pos 6: R-no change
FILTER Pos 7: NB8560-no change
GRATING2: Mirror-no change
GRATING3: 830G
GRATING4: 1200G-no change

No changes required for any
position in the following:
2017-05-24 08:05
 00:00 
SupportDome TempsAfter 10:00 AM, Set both Dome temps to the MKWS forecasted 8:00 PM temperatures, found here: http://hokukea.soest.hawaii.edu/forecast/mko/ Please log temps in comments.
2016-02-07 14:02
 00:00 
SupportHIRESCross-Disperser Change:

Please remember to wear a mask.
Do ***NOT*** calibrate after installation.
Please send an email to hires_info when completed.
Remove: Blue
Install: Red
2017-05-23 16:05
   
SupportDEIMOSFill LN in afternoon
2007-01-11 14:01
   
SupportK1RELK1 Release Mode
2016-07-15 08:07
   
SupportK2RELK2 Release Mode
2016-07-15 08:07
   
SupportLRIS-ADCK1 Instrument & Guider NOT READY
2007-01-11 14:01
   
SupportDEIMOSK2 Instrument & Guider NOT READY
2007-01-11 14:01
   
SupportK1K1 Telescope NOT READY
2007-04-30 17:04
   
SupportK2K2 Telescope NOT READY
2007-01-11 14:01
   
SupportLRISFill LN after telescope initialization
2007-01-11 14:01
   
SupportESIFill LN after telescope initialization
2009-06-04 12:06
    

The server is: mysqlserver

The datebook is: datebook

SIAS version 4.24